Circular No. 7764

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE Further to the supernovae reported on IAUC 7763 by the Supernova

Cosmology Project’s 2001 search (which involved also G. Aldering, R. Amanullah, P. Antilogus, P. Astier, C. Balland, G. Blanc, M. S. Burns, A. Conley, S. E. Deustua, R. Ellis, S. Fabbro, G. Folatelli, A. S. Fruchter, G. Garavini, R. Gibbons, G. Goldhaber, A. Goobar, D. E. Groom, D. Hardin, I. Hook, D. A. Howell, M. Irwin, A. G. Kim, R. A. Knop, J. M. Levy, C. Lidman, R. McMahon, J. Mendez, M. Mouchet, S. Nobili, P. E. Nugent, R. Pain, N. Panagia, E. Pécontal, C. R. Pennypacker, S. Perlmutter, J. Raux, N. Regnault, P. Ruiz-Lapuente, N. A. Walton, L. Wang, and W. M. Wood-Vasey), most of the objects were discovered around maximum light, and follow-up R and I photometry has been obtained for most of the objects (and infrared J-band photometry for four objects). Spectral types and redshifts (given to 0.001 for host-galaxy redshifts, to 0.01 or more for supernovae redshifts) were obtained at the Keck II (+ LRIS) and Keck I (+ ESI) telescopes and with FORS1 during Apr. 20–26. Apparent supernovae 2001gk, 2001gp, 2001gs, 2001gx, and 2001gz were not visible in the Mar. images to limiting mag I ∼ 25.5.

COMETS C/2001 V3, 2001 V4, 2001 V5, 2001 W3, 2001 W4 (SOHO)

Further to IAUC 7760, D. Hammer reports his measurements of additional Kreutz sungrazing comets found by A. Mimeev (C/2001 V5), X. Leprette (C/2001 W4), and X. Zhou (remaining comets) on SOHO website images that were taken with the C2 coronagraph. The reduced astrometry and orbital elements by B. G. Marsden appear on MPEC 2001-X02.

Comet  2001 UT  α_{2000}  δ_{2000}
C/2001 V3 Nov. 11.311  15^{h}03.^{m}7  −18^{o}56′
C/2001 V4  12.713  15 09.1  −19 18
C/2001 V5  14.088  15 14.8  −19 41
C/2001 W3  22.329  15 51.0  −21 48
C/2001 W4  26.479  16 09.5  −22 44

COMET C/2000 WM_{1} (LINEAR) Corrigendum. On IAUC 7762, line 3, for images taken with read images

taken by H. Fukushima with

2001 December 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green